Circular No. 8152

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gd IN M74 R. Kotak and W. P. S. Meikle, Imperial College, London, and S. J.

Smartt, Institute of Astronomy, Cambridge, on behalf of the European Supernova Collaboration, report that a spectrum (range ∼ 400–840 nm) of SN 2003gd was obtained at the William Herschel Telescope (+ ISIS spectrograph) by C. Benn on June 14.2 UT. Comparison of a preliminary reduction of the spectrum with spectra of SN 1992H and SN 1987A (Filippenko 1997, ARAA 35, 309) indicates that the new supernova is of type II, roughly 2 months post-explosion. The strong Hα P-Cyg profile extends from −8800 to +7100 km/s in the host-galaxy rest frame.

M. Phillips, Carnegie Observatories, reports that échelle spectroscopy of SN 2003gd, reduced by M. Navarrete from observations obtained on June 14 UT by G. Preston with the Baade 6.5-m telescope at Las Campanas, indicates that this is a type-II event that was discovered ≈ 1 month after explosion.

Further to IAUC 8150, R. H. McNaught provides a better position for SN 2003gd (and revised red mag 13.3 on June 13.84 UT): α = 1^{h}36^{m}42^{s}.65, δ = +15^{o}44′19′′.9 (equinox 2000.0; uncertainties ± 0′′.1 in each coordinate).

S. J. Smartt, J. R. Maund, and M. A. Hendry, Institute of Astronomy, Cambridge, report that the position of SN 2003gd lies on deep pre-explosion archive images from the Hubble Space Telescope (+ WFPC2) and Gemini North (+ GMOS). Differential astrometry, using an image obtained at the William Herschel Telescope (+ Aux Port Camera) by C. R. Benn on June 15.2 UT, indicates the presence of a progenitor star (V = 26.1 ± 0.15) in the 3100-s F606W WFPC2 image at a distance of 0′′.06 ± 0′′.2 from the supernova position. The star has an absolute magnitude of M_{V} = −3.5, assuming the distance and foreground reddening given by Smartt et al. (2002, Ap.J. 572, L147). The apparent progenitor (I ≈ 23.8, V − I = +2.3) is detected on the GMOS i′ band (0′′.6 seeing) but is close to a neighboring star that is resolved in the HST F606W image. The color and absolute magnitude suggest that this star is a red supergiant, probably of spectral type M, with luminosity log (L/L_{⊙}) ≈ 4 and a relatively low mass in the range 8–10 M_{⊙}. Smartt et al. add that they will use the HST override program to observe SN 2003gd (with ACS), so as to obtain a significant reduction in the astrometric error, as well as to determine the extinction in this region. Given the likelihood of a progenitor detection, observations of this supernova at all wavelengths are strongly encouraged.

2003 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green